OpenStudy (anonymous):

Medal A personal trainer has a supply of weight plates to make custom barbells to use for weight-training. The trainer adds 21 1/4 pounds of plates to each side of the dumbbell bar, which itself weighs 7 1/2 pounds. Write a ratio comparing the weight added to each side of the dumbbell bar to the weight of the bar itself, using a fraction in simplest terms.

OpenStudy (anonymous):

oh no it is not!

OpenStudy (anonymous):

adds 21.25 to both sides for a total of \(42.5\) almost missed that part

OpenStudy (anonymous):

put one on top of the other \[\frac{42.5}{7.5}\] for the ratio remove the decimal, reduce the fraction

OpenStudy (anonymous):

56.667

OpenStudy (anonymous):

they want a fraction, not a decimal

OpenStudy (anonymous):

how do i make this as a fraction

OpenStudy (anonymous):

\[\frac{42.5}{7.5}=\frac{425}{75}\] divide top and bottom by \(25\) to reduce to \[\frac{17}{3}\]

OpenStudy (anonymous):

okay so the answer is 17/3?